OpenStudy (anonymous):

Without plotting any points other than intercepts, draw a possible graph of the following polynomial: f(x) = (x + 8)^3(x + 6)^2(x + 2)(x − 1)3(x − 3)^4(x − 6).

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

is your function like this: \[{\left( {x + 8} \right)^3}{\left( {x + 6} \right)^2}\left( {x + 2} \right){\left( {x - 1} \right)^3}{\left( {x - 3} \right)^4}\left( {x - 6} \right)\]

OpenStudy (anonymous):

Yes :)

OpenStudy (michele_laino):

the graph of your function crosses the x-axis at points x=-8, x=-6, x=-2, x=1, x=3, and x=6 since at those points your function: \[f\left( x \right) = {\left( {x + 8} \right)^3}{\left( {x + 6} \right)^2}\left( {x + 2} \right){\left( {x - 1} \right)^3}{\left( {x - 3} \right)^4}\left( {x - 6} \right)\] is equal to zero

OpenStudy (michele_laino):



OpenStudy (michele_laino):

now, we have this: \[f\left( { - 9} \right) = {\left( { - 9 + 8} \right)^3}{\left( { - 9 + 6} \right)^2}\left( { - 9 + 2} \right){\left( { - 9 - 1} \right)^3}{\left( { - 9 - 3} \right)^4}\left( { - 9 - 6} \right) > 0\]

OpenStudy (michele_laino):

whereas: \[f\left( 7 \right) = {\left( {7 + 8} \right)^3}{\left( {7 + 6} \right)^2}\left( {7 + 2} \right){\left( {7 - 1} \right)^3}{\left( {7 - 3} \right)^4}\left( {7 - 6} \right) > 0\]

OpenStudy (michele_laino):

so the graph of your function can be like this: